OpenStudy (anonymous):

A particle of mass m fall freely from rest under gravity for a time 1/2k.how find the distance it has fallen and the speed attained.

OpenStudy (ash2326):

s=ut+1/2 at^2 u is the initial velocity t a= acceleration here it's gravity t=time s=1/2* 9.8*t^2 substitute t to find the distance covered

OpenStudy (ash2326):

for speed v=u+at u=0 v is the final velocity v=9.8*t substitute t to find the velocity

OpenStudy (anonymous):

s=g/8k2 and v=g/2k isn't?

OpenStudy (ash2326):

yeah